OpenStudy (4meisu):

Given that 2sin^2 θ + sinθ - 1 = 0, find the two values for sin θ

OpenStudy (anonymous):

2sin² θ + sin θ - 1 = 0 (2sin θ - 1)(sin θ + 1) = 0

OpenStudy (lgbasallote):

imagine \(\sin \theta = x\) so this thingy will become 2x^2 + x - 1 = 0 do you know how to solve for x there?

OpenStudy (anonymous):

so much can you understand ?

OpenStudy (anonymous):

@4meisu ?

OpenStudy (4meisu):

No I do not know how to solve for x.. What's next?

OpenStudy (anonymous):

Now we can split that into two equations 2sin θ - 1 = 0 and sin θ + 1 = 0 this much understood ?

OpenStudy (4meisu):

yes

OpenStudy (4meisu):

sinθ = 1/2 and sinθ = -1 ?

OpenStudy (anonymous):

Solving the first one: 2sin θ - 1 = 0 sin θ = 1/2 θ = 30°, 150°

OpenStudy (anonymous):

Second: sin θ + 1 = 0 sin θ = -1 θ = 270°

OpenStudy (4meisu):

Okay thank you

OpenStudy (anonymous):

:)

OpenStudy (4meisu):

The question is asking for only two values for theta though, which one would we choose?